the area of the triangle formed by the tangent at the pt. (a,b) to the circle x2 +y2 =r2and the coordinate axis


 

Grade:12

1 Answers

Jit Mitra
12 years ago

Equation of tangent is 

 

ax+by-r^2 = 0

In x-y intercept form

 

x/(r^2/a) + y/(r^2/b) = 1

 

x-intercept = r^2/b

y-intercept = r^2/b

 

Area of triangle = 1/2*r^2/b*r^2/b = r4/(2ab)